# Training-Free Token Pruning via Zeroth-Order Gradient Estimation in Vision-Language Models

### v1

@@ 3.3 ZOO-Prune: ZerOth-Order gradient estimation for token pruning
4 Experiments
[…]
Table 2: Performance Comparison on LLaVA-1.5-13B.
Method
GQA
Acc. $\uparrow$
MMB
Acc. $\uparrow$
MME
P+C $\uparrow$
POPE
F1 $\uparrow$
SQA
Acc. $\uparrow$
VQAV2
Acc. $\uparrow$
VQAText
Acc. $\uparrow$
MMMU
Acc. $\uparrow$
SEED-I
Acc. $\uparrow$
Avg.
$\uparrow$
Total 576 Tokens
100% 100% 100% 100% 100% 100% 100% 100% 100%
Retain 192 Tokens $\downarrow$ 66.7%
Retain 128 Tokens $\downarrow$ 77.8%
Retain 64 Tokens $\downarrow$ 88.9%
@@ 4.1 Comparison on Diverse Tasks
Results on LLaVA-1.5-7B and 13B.
As shown in Tables 1 and 2, ZOO-Prune consistently outperforms the state-of-the-art training-free pruning methods, especially under aggressive compression. On LLaVA-1.5-7B, it preserves 95.5% performance with only 64 tokens, surpassing DivPrune (Alvar et al., 2025) (94.8%) and far exceeding attention-based methods such as FastV (Chen et al., 2024a), which retains three times more tokens but drops to 89.6%. On LLaVA-1.5-13B, ZOO-Prune achieves 96.5% with 64 tokens and outperforms DivPrune by 1.7%, confirming that combining sensitivity and diversity effectively preserves reasoning under extreme pruning.
@@
We further evaluate ZOO-Prune on LLaVA-NeXT-7B, which processes up to 2880 tokens. As reported in Table 3, ZOO-Prune maintains 98.3% performance when pruning 77.8% of tokens, outperforming VisionZip (97.5%) and nearly matching the baseline. Even with a 94.4% pruning (160 tokens), it can still maintain a 95.4% performance, demonstrating the scalability of sensitivity-aware diversity selection under high-resolution compression.
@@
To validate the generalizability of ZOO-Prune, we evaluated it on Qwen2.5-VL-7B, a distinct VLM variant beyond the LLaVA family with dynamic-resolution inputs. As shown in Table 4, ZOO-Prune continues to lead, achieving 96.2% performance with a 20% token budget and 90.8% with only 10% of the tokens, consistently surpassing both VisionZip and DivPrune. These results confirm that ZOO-Prune generalizes robustly across diverse VLM architectures, effectively preserving task performance even with long, variable-length token sequences.
Figure 4: Hyperparameter sensitivity on POPE with LLaVA-1.5-7B: (a) effect of step size $h$, (b) effect of the number of perturbation directions $m$.
Table 4: Performance comparison on Qwen2.5-VL-7B with a dynamic resolution setting.
Method GQA MMB MME POPE Avg.
Baseline (Full Tokens)
Qwen2.5-VL-7B 60.84 84.10 2310 86.3 100%
Retain 20% Tokens
VisionZip (CVPR 2025) 57.27 79.72 2221 83.89 95.6%
DivPrune (CVPR 2025) 60.05 79.55 2173 83.42 96.0%
ZOO-Prune (Ours) 58.81 80.60 2201 84.17 96.2%
Retain 10% Tokens
VisionZip (CVPR 2025) 54.09 76.03 1937 78.97 88.7%
DivPrune (CVPR 2025) 55.49 76.03 2054 79.05 90.5%
ZOO-Prune (Ours) 55.45 76.28 2018 80.99 90.8%
4.2 Ablation Studies and Analysis
@@
We ablate four variants on LLaVA-NeXT-7B (Table 5): Sensitivity-only, Diversity-only (DivPrune), and two Fusion strategies (sum, multiply). Sensitivity-only selects task-critical tokens and excels on reasoning, but lags on TextVQA where context is vital. Diversity-only offers broad coverage yet misses key cues. Combining both is consistently better: Fusion (Multiply) delivers 98.3% at 22.2% tokens with no extra hyperparameters. These results show that sensitivity and diversity are complementary and jointly essential for training-free token pruning.
Hyperparameter Analysis.
We analyze the sensitivity of ZOO-Prune to its two hyperparameters $m$ (number of perturbations) and $h$ (step size) on the POPE benchmark (Fig. 4). Performance is consistently stable across a wide range: accuracy remains highly consistent as $m$ varies from 16 to 160 and is similarly insensitive to $h$ over $1\text{e-}4,1]$. We adopt $m=64$ to reduce variance at negligible cost, and fix $h=0.01$ for all experiments, thereby eliminating the need for task-specific tuning.
@@
Figure 5: Inference efficiency on the ScienceQA benchmark relative to the LLaVA-NeXT-7B baseline. The left panel reports prefilling time and the right panel shows end-to-end (E2E) latency. Speedup factors over the baseline are annotated inside each bar.
@@
Visualization.
Fig. 6 qualitatively compares ZOO-Prune with DivPrune under aggressive pruning on GQA. Although DivPrune preserves feature diversity, it often drops critical tokens, for example, omitting most dog tokens and predicting “bear,” or focusing on a window instead of the queried object. In contrast, ZOO-Prune uses zeroth-order sensitivity to preserve key entities like the dog and bench. By combining sensitivity and diversity, it maintains correct predictions even at a 64-token budget, underscoring that sensitivity-aware diversity is essential for semantic integrity and reliable reasoning under heavy compression. See Appendix E for more examples.
Figure 6: Qualitative comparison of pruned token masks from DivPrune and ZOO-Prune on the GQA benchmark under varying token budgets.
@@ 5 Conclusion
In this paper, we presented ZOO-Prune, a training-free and attention-free token pruning framework for vision–language models that unifies zeroth-order sensitivity estimation with diversity-aware selection. By leveraging lightweight projection-layer responses, our approach efficiently measures token importance without backpropagation or additional end-to-end passes, while Sensitivity-Aware Diversity Selection ensures both informativeness and representational coverage. Extensive experiments across multiple benchmarks show that ZOO-Prune achieves state-of-the-art performance among training-free pruning methods, enabling substantial reductions in inference cost with minimal loss in accuracy. We believe this work highlights the potential of gradient-free sensitivity analysis for efficient multimodal learning and opens promising directions for scalable model compression.
[…]
Reproducibility Statement.
Our method is detailed in Section 3 and Algorithm 1, with theoretical analysis in Appendix B.1. All experimental setup details, including models, benchmarks, and hyperparameters, are provided in Appendix C. Our work, built upon the open-source lmms-eval framework, will be released with full source code to ensure reproducibility.
@@
This appendix supplements the main paper with the theoretical analysis of Proposition 3.1, additional details on the experimental configuration and evaluation strategy, and further qualitative results that illustrate the token preservation patterns across a range of scenarios. The contents are organized as follows:
  • Appendix A: Implementation details for KDE experiments in Section 3.2.
@@
  • Appendix D: Further Inference Efficiency Analysis: FLOPs.
@@
  • Appendix F: The Use of Large Language Models (LLMs).
Appendix A Experimental Details for Spearman Correlation Analysis
For the correlation analysis in Fig. 2, we measured the agreement between token-importance rankings derived from the vision encoder and the projection layer. Specifically, we selected 50 random samples per dataset from MMMU and POPE. For each sample, token sensitivities were first computed using RGE at both the vision encoder output and the projection layer. To ensure stable ranking comparisons, we applied a threshold of 0.5 to filter out low-sensitivity tokens before computing ranks. The Spearman’s rank correlation coefficient was then calculated for each sample, and the distribution across 50 samples was visualized using a kernel density estimate (KDE) plot.
The resulting average Spearman correlations were 0.55 for MMMU and 0.49 for POPE, indicating a consistent alignment between token rankings obtained at the projection layer and those from the full vision encoder. This confirms that the projection layer can serve as a reliable proxy for token-level importance estimation while significantly reducing computational overhead.
@@ B.1 Propositional Proof
Let $M:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be differentiable at $x\in\mathbb{R}^{n}$ with Jacobian $J(x)=\nabla M(x)$. Let $u\sim\mathcal{N}(0,I_{n})$ be an isotropic Gaussian perturbation and $h>0$ a small step size. Define the finite-difference sensitivity $S(x)=\mathbb{E}_{u}\left[\left\|\frac{M(x+hu)-M(x-hu)}{2h}\right\|_{2}\right].$ Then, for sufficiently small $h$,
$S(x)=\mathbb{E}_{u}\big[\|J(x)u\|_{2}\big]+O(h^{2}).$
@@
$\left\|\frac{M(x+hu)-M(x-hu)}{2h}\right\|_{2}=\|J(x)u+O(h^{2})\|_{2}=\|J(x)u\|_{2}+O(h^{2}).$
@@ C.1 Model Settings
We evaluate the effectiveness of ZOO-Prune on widely used VLMs, including LLaVA-v1.5-7B (Liu et al., 2024a) [^1] https://huggingface.co/liuhaotian/llava-v1.5-7B, LLaVA-v1.5-13B (Liu et al., 2024a) [^2] https://huggingface.co/liuhaotian/llava-v1.5-13B, and LLaVA-1.6-7B (Liu et al., 2024b) [^3] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b (also referred to as LLaVA-NeXT-7B), and Qwen2.5-VL-7B (Bai et al., 2025) [^4] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct-AWQ. All LLaVA models adopt the CLIP (Radford et al., 2021) as the vision encoder and Vicuna (Chiang et al., 2023) as the base language model.
@@ C.2 Implementation Details
For pruning ratios, we adopt 66.7%/77.8%/88.9% for LLaVA-v1.5 and 77.8%/88.9%/94.4% for LLaVA-NeXT-7B. In the latter case, we follow the implementation of VisionZip (Yang et al., 2025), where the model dynamically samples up to five image patches, resulting in as many as 2880 vision tokens. For example, with a pruning budget of 160 tokens, we retain 32 tokens per patch across five patches (32 × 5 = 160). If fewer patches are sampled (e.g., four), the number of retained tokens is adjusted proportionally (e.g., 128 tokens for 160/2880). We applied a low-rank factorization ($k=128$) to the MM-projector layers to further boost efficiency on LLaVA-Next, due to the large number of visual tokens. For the dynamic-resolution Qwen2.5-VL-7B, we evaluate at 10% and 20% token retention rates.
Finally, as validated in ablation, our method remains robust across different hyperparameter choices. Unless otherwise noted, we fix the perturbation hyperparameters to $m=64$ and $h=0.01$ for all experiments. Evaluation is performed using the lmms-eval [^5] https://github.com/EvolvingLMMs-Lab/lmms-eval framework under official protocols and metrics.
Table A: Summary of primary evaluation metrics for each benchmark.

### v2

@@ 4 Experiments
ToMe (ICLR 2023) 54.30 60.50 1563.00 72.40 65.20 68.00 52.10 88.49%
FastV (ECCV 2024) 52.70 61.20 1612.00 64.80 67.30 67.10 52.50 87.75%
HiRED (AAAI 2025) 58.70 62.80 1737.00 82.80 68.40 74.90 47.40 93.84%
VisionZip (CVPR 2025) 59.30 63.00 1782.60 85.30 68.90 76.80 57.30 97.66%
DivPrune (CVPR 2025) 59.97 62.54 1762.23 87.00 68.91 76.87 56.97 97.78%
PyramidDrop (CVPR 2025) 57.10 63.20 1766.00 82.30 68.80 75.10 56.10 95.95%
SparseVLM (ICML 2025) 57.60 62.50 1721.00 83.60 69.10 75.60 56.10 95.93%
ZOO-Prune (Ours) 60.03 62.89 1781.66 87.24 69.16 77.34 57.30 98.27%
@@
ToMe (ICLR 2023) 52.40 53.30 1343.00 62.80 59.60 63.00 49.10 80.38%
FastV (ECCV 2024) 49.60 56.10 1490.00 59.60 60.20 61.80 50.60 81.22%
HiRED (AAAI 2025) 57.20 61.50 1710.00 79.80 68.10 73.40 46.10 91.84%
VisionZip (CVPR 2025) 57.60 62.00 1761.70 83.20 68.90 75.60 56.80 96.20%
DivPrune (CVPR 2025) 59.25 62.03 1718.22 86.72 68.96 75.96 56.06 96.73%
PyramidDrop (CVPR 2025) 56.00 61.00 1644.00 82.30 68.30 72.90 55.10 93.52%
SparseVLM (ICML 2025) 56.00 60.00 1696.00 80.50 67.10 73.80 54.90 93.27%
ZOO-Prune (Ours) 59.49 61.86 1751.60 87.13 68.91 76.57 57.87 97.62%
@@
ToMe (ICLR 2023) 48.60 43.70 1138.00 52.50 50.00 57.10 45.30 70.12%
FastV (ECCV 2024) 46.10 48.00 1256.00 48.00 51.10 55.00 47.80 71.10%
HiRED (AAAI 2025) 54.60 60.20 1599.00 73.60 68.20 69.70 44.20 87.95%
VisionZip (CVPR 2025) 55.10 60.10 1690.00 77.00 69.00 72.40 55.50 92.74%
DivPrune (CVPR 2025) 57.78 59.28 1674.40 85.56 68.17 74.11 54.69 94.42%
PyramidDrop (CVPR 2025) 41.90 33.30 1092.00 55.90 68.60 69.20 45.90 72.66%
SparseVLM (ICML 2025) 52.70 56.20 1505.00 75.10 62.20 68.20 51.80 86.52%
ZOO-Prune (Ours) 58.47 60.22 1675.59 85.86 68.27 75.02 55.35 95.20%
@@ 4.1 Comparison on Diverse Tasks
Results on LLaVA-1.5-7B.
As shown in Table 1, ZOO-Prune consistently outperforms the state-of-the-art training-free pruning methods, especially under aggressive compression. On LLaVA-1.5-7B, it preserves 95.20% performance with only 64 tokens, surpassing DivPrune [1] (94.40%) and far exceeding attention-based methods such as FastV [10], which retains three times more tokens but drops to 87.75%. Additional results on LLaVA-1.5-13B in Appendix D.
Table 2: Performance comparison on Qwen2.5-VL-7B.
Method GQA MMB MME POPE Avg.
Baseline (Full Tokens)
Qwen2.5-VL-7B 60.84 84.10 2310 86.3 100%
Retain 20% Tokens
VisionZip (CVPR 2025) 57.27 79.72 2221 83.89 95.6%
DivPrune (CVPR 2025) 60.05 79.55 2173 83.42 96.0%
ZOO-Prune (Ours) 58.81 80.60 2201 84.17 96.2%
Retain 10% Tokens
VisionZip (CVPR 2025) 54.09 76.03 1937 78.97 88.7%
DivPrune (CVPR 2025) 55.49 76.03 2054 79.05 90.5%
ZOO-Prune (Ours) 55.45 76.28 2018 80.99 90.8%
Figure 4: Hyperparameter sensitivity on POPE with LLaVA-1.5-7B: (a) small step size $h$, (b) number of perturbation directions $m$.
@@
We further evaluate ZOO-Prune on LLaVA-NeXT-7B, which processes up to 2880 tokens. As reported in Table 3, ZOO-Prune maintains 98.3% performance when pruning 77.8% of tokens, outperforming VisionZip and nearly matching the baseline. Even with a 94.4% pruning (160 tokens), it can maintain a 95.4% performance, demonstrating the scalability of sensitivity-aware diversity selection under high-resolution compression.
@@
To validate the generalizability of ZOO-Prune, we tested it on Qwen2.5-VL-7B, a distinct VLM variant beyond LLaVA family with dynamic-resolution inputs. As shown in Table 2, ZOO-Prune continues to lead, achieving 96.2% performance with a 20% token budget and 90.8% with only 10% of the tokens, consistently surpassing both VisionZip and DivPrune. These results confirm that ZOO-Prune generalizes robustly across diverse VLM architectures, effectively preserving task performance even with long, variable-length token sequences.
4.2 Ablation Studies and Further Analysis
@@
We ablate four variants on LLaVA-NeXT-7B (Table 4): Sensitivity-only, Diversity-only (DivPrune), and two Fusion strategies (sum, multiply). Sensitivity-only selects task-critical tokens and excels on reasoning, but lags on TextVQA where context is vital. Diversity-only offers broad coverage yet misses key cues. Combining both is consistently better: Fusion (Multiply) delivers 98.3% at 22.2% tokens without extra hyperparameters. It means that sensitivity and diversity are complementary and jointly essential for training-free token pruning.
Figure 5: Qualitative comparison on the GQA benchmark. Attention-based methods suffer from positional bias or redundant token clusters. Diversity-based pruning spreads tokens broadly but lacks semantic focus. The ZOO-based sensitivity captures output-related tokens but overlooks spatial coverage. Our ZOO-Prune jointly optimizes sensitivity and diversity for balanced selection across compression ratios.
Figure 6: ZOO-based Sensitivity vs. Attention-guided Pruning on POPE. Our ZOO-based sensitivity (Sens) metric consistently outperforms both text-visual (T2V Attn) and visual-visual (V2V Attn) attention-based pruning, even when combined with diversity (Div).
@@
Figure 7: Inference efficiency on the POPE benchmark relative to the LLaVA-NeXT-7B baseline. The left scatter plot reports prefilling time, and the right scatter plot shows end-to-end (E2E) latency. Each point represents a method–token-count pair, illustrating the trade-off between computation cost and F1 score performance.
Hyperparameter Analysis.
We analyze the sensitivity of ZOO-Prune to its two hyperparameters $m$ (number of perturbations) and $h$ (step size) on the POPE benchmark (Fig. 4). Performance is stable across a wide range: accuracy remains consistent for m=16m=16160160 and is similarly insensitive to $h$ over $\text{e-}411$. We adopt $m=64$ to reduce variance at negligible cost, and fix $h=0.01$ for all experiments, eliminating the need for task-specific tuning.
Revisiting Token Importance: Sensitivity vs. Attention vs. Diversity.
We compare our zeroth-order sensitivity metric against two main categories of training-free baselines: attention-based and diversity-based. For the attention-based methods, we implement two variants: (1) a text-visual (T2V) approach using attention scores after layer 3, following FastV [10], and (2) a visual-visual (V2V) approach like VisionZip [42] that averages token attention scores. For the diversity-based pruning, we implement the max-min feature diversity selection from DivPrune [1].
As shown in Fig. 5, attention-guided pruning suffers from inherently limitations: T2V attention exhibits positional bias by overweighting tokens physically near the query (often in lower image regions), while V2V attention preserves numerous duplicate tokens. These issues make attention-based pruning unstable under aggressive compression. Diversity-only pruning improves spatial coverage but treats tokens uniformly, often discarding semantically important regions. Our ZOO-based sensitivity quantifies each token’s impact on the model output, providing a stable and architecture-agnostic importance signal.
We also evaluate the behavioral and consistency aspects of each method through qualitative analyses. As shown in Fig. 7, sensitivity-based pruning consistently outperforms attention-driven variants across all compression ratios. When combined with diversity in ZOO-Prune, it retains high-sensitivity tokens while ensuring sufficient content coverage, achieving a markedly superior accuracy–efficiency trade-off. Overall, these quantitative and qualitative results show that ZOO-Prune unifies sensitivity and diversity into a robust and broadly applicable token pruning framework. See more in Appendix E.
@@ 5 Conclusion
In this paper, we presented ZOO-Prune, a training-free and attention-free token pruning framework for vision–language models that unifies zeroth-order sensitivity estimation with diversity-aware selection. By leveraging projection-layer responses, our method measures token importance without backpropagation, while Sensitivity-Aware Diversity Selection preserves both informativeness and coverage. Extensive experiments across multiple benchmarks show that ZOO-Prune achieves state-of-the-art accuracy among training-free pruning methods and reduces inference cost with minimal loss. In addition, the method remains stable across architectures and pruning ratios, demonstrating strong generality. Interestingly, even at extremely high pruning levels, the sensitivity-based ranking reliably preserved the same set of task-critical regions across models. We believe these results highlight the practical value of gradient-free sensitivity signals for efficient multimodal learning and point toward broader applications in scalable token and feature compression.
@@
This appendix supplements the main paper with the theoretical analysis of Proposition 3.1, additional details on the experimental configuration and evaluation strategy, and further quantitative and qualitative results that illustrate the token preservation patterns across a range of scenarios. The contents are organized as follows:
  • Appendix A: Implementation details for KDE experiments in Section 3.2 of the main paper.
@@
  • Appendix D: Further quantitative results and inference efficiency analysis.
@@
  • Appendix F: Discussion and Future Work.
Appendix A Spearman Correlation Setup
For the correlation analysis in Fig. 2 of the main paper, we examined how closely the token-importance rankings from the vision encoder matched those from the projection layer. Specifically, we selected 50 random samples per dataset from MMMU and POPE. For each sample, token sensitivities were first computed using RGE at both the vision encoder output and the projection layer. To ensure stable ranking comparisons, we applied a 0.5 threshold to filter out low-sensitivity tokens before computing ranks. The Spearman’s rank correlation coefficient was then calculated for each sample, and the distribution across 50 samples was visualized using a kernel density estimate (KDE) plot.
The resulting average Spearman correlations were 0.55 for MMMU and 0.49 for POPE, indicating a consistent alignment between token rankings obtained at the projection layer and those from the full vision encoder. This confirms that the projection layer can serve as a reliable proxy for token-level importance estimation while significantly reducing the computational overhead.
@@ B.1 Proof of Proposition
Let $M:\!\mathbb{R}^{n}\!\to\!\mathbb{R}^{m}$ be differentiable at $x\!\in\!\mathbb{R}^{n}$ with Jacobian $J(x)\!=\!\nabla M(x)$. Let $u\!\sim\!\mathcal{N}(0,I_{n})$ be an isotropic Gaussian perturbation and h>0h\!>\!0 a small step size. Define the finite-difference sensitivity $S(x)\!=\!\mathbb{E}_{u}\left[\left\|\frac{M(x+hu)-M(x-hu)}{2h}\right\|_{2}\right].$ Then, for sufficiently small $h$,
$S(x)=\mathbb{E}_{u}\big[\|J(x)u\|_{2}\big]+O(h^{2}).$ (8)
@@
$\left\|\frac{M(x\!+\!hu)\!-\!M(x\!-\!hu)}{2h}\right\|_{2}$ $=\|J(x)u\!+\!O(h^{2})\|_{2}$ (12)
$=\|J(x)u\|_{2}\!+\!O(h^{2}).$ (13)
@@ C.1 Model Settings
We evaluate the effectiveness of ZOO-Prune on widely used VLMs, including LLaVA-v1.5-7B [24], LLaVA-v1.5-13B [24], and LLaVA-1.6-7B [25] (also referred to as LLaVA-NeXT-7B), and Qwen2.5-VL-7B [4]. All LLaVA models adopt the CLIP [32] as the vision encoder and Vicuna [14] as the base language model.
@@ C.2 Implementation Details
For pruning ratios, we adopt 66.7% / 77.8% / 88.9% for LLaVA-v1.5 and 77.8% / 88.9% / 94.4% for LLaVA-NeXT-7B. In the latter case, we follow the implementation of VisionZip [42], where the model dynamically samples up to five image patches, resulting in as many as 2,880 vision tokens. For example, with a pruning budget of 160 tokens, we retain 32 tokens per patch across five patches (32 × 5 = 160). If fewer patches are sampled (e.g., four), the number of retained tokens is adjusted proportionally (e.g., 128 tokens for 160 / 2880). We applied a low-rank factorization ($k=128$) to the MM-projector layers to further boost efficiency on LLaVA-NeXT, due to the large number of visual tokens. For the dynamic-resolution Qwen2.5-VL-7B, we evaluate at 10% and 20% token retention rates.
Finally, as validated in ablation, our method remains robust across different hyperparameter choices. Unless otherwise noted, we fix the perturbation hyperparameters to $m=64$ and $h=0.01$ for all experiments. Evaluation is performed using the lmms-eval [47] framework under official protocols and metrics.
Table A: Summary of primary evaluation metrics.
@@
Table B: Performance Comparison on LLaVA-1.5-7B.
Method
GQA
Acc. $\uparrow$
MMB
Acc. $\uparrow$
MME
P+C $\uparrow$
POPE
F1 $\uparrow$
SQA
Acc. $\uparrow$
VQAV2
Acc. $\uparrow$
VQAText
Acc. $\uparrow$
MMMU
Acc. $\uparrow$
SEED
Acc. $\uparrow$
Avg.
$\uparrow$
Total 576 Tokens
Retain 192 Tokens $\downarrow$ 66.7%
Retain 128 Tokens $\downarrow$ 77.8%
Retain 64 Tokens $\downarrow$ 88.9%